Tom and Sally travel to Chicago. Sally drives the first part of the trip averaging 64 mph. Tom drives the rest of the way averaging 70mph. As it works out, Tom drives one-half hour longer than Sally. If the total distance of the trip was 343.2 miles, for how many hours did Tom drive? I need a chart of the person, Rate, Time, and Distance.
Answer by stanbon(75887)

Tom and Sally travel to Chicago. Sally drives the first part of the trip averaging 64 mph. Tom drives the rest of the way averaging 70mph. As it works out, Tom drives one-half hour longer than Sally. If the total distance of the trip was 343.2 miles, for how many hours did Tom drive? I need a chart of the person, Rate, Time, and Distance.
----
Sally DATA:
rate = 64 mph ; time = x hrs ; distance = r*t = 64x miles
-----------
Tom DATA:
rate = 70 mph ; time = x+(1/2) hrs ; distance = r*t = 70x + 35 miles
----------------
Equation:
dist + dist = 343.2 miles
64x + 70x + 35 = 343.2
---
134x = 308.2
x = 2.3 hrs (Sally's time)
----
x+0.5 = 2.8 hrs = 2hr 48 min (Tom's time)
--------------------
Cheers,
Stan H.
------------